OpenStudy (laurenarg):

A circle has a diameter with endpoints (-8, 2) and (-2, 6). What is the equation of the circle? r2 = (x - 3)2 + (y + 4)2 r2 = (x - 5)2 + (y + 4)2 r2 = (x + 5)2 + (y - 4)2 r2 = (x + 3)2 + (y - 4)2

OpenStudy (phi):

I would find the average of the x values of (-8, 2) and (-2, 6) and the average of the y values of those two points. that will be the center of the circle if you get (h,k) the equation would be (x-h)^2 + (y-k)^2 = r^2

OpenStudy (laurenarg):

What is (h.k).. I neveer quite understood that...

TheSmartOne (thesmartone):

(h, k) is the center, but I'll let Phi take over :p

OpenStudy (phi):

(h,k) is just to show a pair of numbers. we use (h,k) as the (x,y) values of the center point of the circle. but before worrying about that, can you find the average of -8 and -2 ?

OpenStudy (phi):

(to find the average , add up the two numbers, and then divide by 2)

OpenStudy (laurenarg):

yeah I have x=-5 and y= 4

OpenStudy (phi):

now you use those numbers in the "template" (x-h)^2 + (y-k)^2 = r^2 where (h,k) is your (-5,4)

OpenStudy (laurenarg):

(x+5)^2 + (y-4)^2=r^2

OpenStudy (phi):

yes, and if we wanted we could find the radius using the distance formula between the center and one of the points they gave us. I think we get r^2 = 13

OpenStudy (laurenarg):

Ok thank you!! Could you help with some more please...

OpenStudy (phi):

we could use for example \[ ( x_0, y_0) \] instead of (h,k) to show a pair of numbers but it's harder to type

OpenStudy (laurenarg):

The vertices of an isosceles triangle are A(3, 6), B(7, 2), and C(4, 3). What is the equation of the triangle's line of symmetry? y = x + 1 y = x - 1 y = x y = -x + 1

OpenStudy (phi):

I would plot the 3 points first, to see what is going on.

OpenStudy (phi):

can you sketch the plot ?

OpenStudy (laurenarg):

Yeah, hold on

OpenStudy (laurenarg):

Alright..now what?

OpenStudy (phi):

the "axis of symmetry" is the line where you can fold the triangle in half, and the two sides lay on top of each other.

OpenStudy (phi):

any idea which point this "axis" goes through ?

OpenStudy (laurenarg):

Point C is the middle point of the angle...

OpenStudy (phi):

yes. any idea what angle this "axis" makes with the "base" side i.e line segment AB ?

OpenStudy (laurenarg):



OpenStudy (laurenarg):

An obtuse angle...?

OpenStudy (phi):



OpenStudy (phi):

the axis is the same as the line that bisects the base, and forms a right angle with the base. It's easier to see if you rotate the triangle so it looks "normal" any way, we can find the slope of the axis, because it is perpendicular to line AB if line AB has slope m, the axis will have slope -1/m (negative reciprocal) I would find the slope of AB flip it and multiply by -1 that will be the slope of the axis then use y = mx + b put in the point (4,3) and solve for b

OpenStudy (laurenarg):

I don't understand

OpenStudy (phi):

can you find the slope of AB ?

OpenStudy (laurenarg):

Yes

OpenStudy (phi):

what do you get ?

OpenStudy (laurenarg):

1

OpenStudy (phi):

close. but that is not the slope for line AB change in y divided by change in x

OpenStudy (phi):

can you post what you did?

OpenStudy (laurenarg):

-1 Sorry wasn't paying attention

OpenStudy (phi):

yes, that is better. next, we use this fact if a line has a slope of m a line perpendicular to that line will have slope -1/m (it's one of those facts they expect you to know... to answer this question) so m= -1 write -1 divided by m: -1/-1 that simplifies to 1 1 is the slope of axis line.

OpenStudy (phi):

if you look at your choices, we can cross of the last choice. it has the wrong slope we have a point (4,3) and a slope m=1 we need to find the equation of a line. Do you know how to do that ?

OpenStudy (laurenarg):

No...

OpenStudy (phi):

one way is use y = m x + b we know m is 1 so we can write y = 1x + b or just y = x+b this equation has to work for the point (4,3) (point C) put in 4 for x and 3 for y: 3= 4 + b now solve for b can you do that ?

OpenStudy (phi):

I would add -4 to both sides, and simplify

OpenStudy (laurenarg):

b=-1

OpenStudy (phi):

so m=1 and b=-1 for the axis line: y = 1x + -1 or y = x -1 if we pretty it up a little.